OpenStudy (anonymous):

Im trying to find the tangent ratio of an angle in a triangle. I think that you're suppose to use the converse of tan when doing this. I know that the opposite side from angle A is 6, and the adjacent side is 8. Therefore, tan^-1(A)=6/8. Please help me simplify this!

OpenStudy (anonymous):

Hi Welcome to Openstudy! If the opposite is 6 and the adjacent is 8, then the tangent ratio will be \(\sf \large tan (A) = \frac{6}{8}\) if you're trying to find the measure of the angle A, then you have to use the inverse of tan, but it will look like this: \(\sf \large (A) =tan^{-1} (\frac{6}{8})\) not \(\sf \large tan^{-1} (A) =tan^{-1} (\frac{6}{8})\) hope this helps :) if you have more questions don't hesitate to ask

OpenStudy (anonymous):

Thank you so much! <3

OpenStudy (anonymous):

you're always welcome! ^_^